OpenStudy (anonymous):

Please help me with this.. When given the vertex and the y-intercept of a parabola, use this vertex form of the equation to find the value for a and the final equation : (y-k)=a(x-h)^2 Find the equation of the parabola with vertex (1,-108) and y-intercept (0,-5). Write your answer in vertex form. And PLEASE SHOW ALL YOUR WORK.

OpenStudy (anonymous):

Plssss help

Directrix (directrix):

@AlexisDeniseBurns You'll need to come back online to help with this.

OpenStudy (anonymous):

I'm online

Directrix (directrix):

Okay. We will get started.

Directrix (directrix):

(y-k)=a(x-h)^2 (h,k) in the formula is for the vertex of the parabola. The vertex is given as vertex (1,-108)

Directrix (directrix):

In this equation, substitute: 1 for h -108 for k and post what you get when you do that, okay?

OpenStudy (anonymous):

(y--108)=a(x-1)^2

Directrix (directrix):

Which gives: (y+108)=a * (x-1)^2

OpenStudy (anonymous):

Yeah I understand

Directrix (directrix):

The parabola has y-intercept (0,-5). The y-intercept is a point on the parabola.

Directrix (directrix):

Take this (y+108)=a * (x-1)^2 and replace x by 0 and y by -5 Post what you get when you do that, okay?

OpenStudy (anonymous):

(-5+108)=a * (0-1)^2

Directrix (directrix):

(-5+108)=a * (0-1)^2 103 = a* (-1)^2 103 = a*1 103 = a Do you agree?

OpenStudy (anonymous):

I agree

Directrix (directrix):

The last task is to write the equation of the parabola in vertex form.

Directrix (directrix):

Go back up the thread to the equation you wrote with the coordinates of the vertex. (y+108)=a * (x-1)^2

Directrix (directrix):

We know now that a = 103

Directrix (directrix):

So, what is the final answer for the equation of the parabola in vertex form?

Directrix (directrix):

Take this and replace a with 103. (y+108)=a * (x-1)^2

OpenStudy (anonymous):

(y+108)=103*(x-1)^2

OpenStudy (anonymous):

Is my final equation (-5+108)=103(0-1)^2

Directrix (directrix):

The vertex form for a parabola is this: y=a(x-h)^2 + k. Your question gives this as the vertex form of a parabola: (y-k)=a(x-h)^2 The answer will be this according to your question: (y+108)=103*(x-1)^2 Or, according to the definition of vertex form given elsewhere, y=103*(x-1)^2 - 108

Directrix (directrix):

I would go with this because it is in what I learned as the vertex form of a parabola: y=103*(x-1)^2 - 108

OpenStudy (anonymous):

Which one is my final equation and which one Is the equation of the parabola

Directrix (directrix):

But, look in your text to see what is given to be the vertex form of a parabola. From your question, it appears that this would be the answer for this question: (y+108)=103*(x-1)^2

Directrix (directrix):

>>Is my final equation (-5+108)=103(0-1)^2 No.

OpenStudy (anonymous):

What exactly should I write on my paper, that would answer all that the question is asking

Directrix (directrix):

(y-k)=a(x-h)^2 (h,k) in the formula is for the vertex of the parabola. The vertex is given as vertex (1,-108) (y+108)=a * (x-1)^2 The parabola has y-intercept (0,-5). The y-intercept is a point on the parabola. (-5+108)=a * (0-1)^2 103 = a* (-1)^2 (y+108)=103*(x-1)^2 103 = a*1 103 = a

OpenStudy (anonymous):

Thank u SOOOOOOOOOOOOOO much

Directrix (directrix):

You are welcome.